OpenStudy (s):

Find the derivative and simplify : y=(3x^2+5x+1)/(3-x^2) ?

OpenStudy (mertsj):

\[y'=\frac{(3-x^2)(6x+5)-(3x^2+5x+1)(-2x)}{(3-x^2)^2}\]

OpenStudy (mertsj):

Probably want to do some algebra on that and clean it up a bit.

OpenStudy (s):

I got this derivative, and later i was confused with algebra, and never got the right answer.( should be 5(x+1)(x+3) divided by (3-x^2)^2

OpenStudy (mertsj):

Let's work on the numerator then

OpenStudy (mertsj):

\[(3-x^2)(6x+5)=18x+15-6x^3-5x^2=-6x^3-5x^2+18x+15\]

OpenStudy (mertsj):

Set that aside for a minute while we get \[-2x(3x^2+5x+1)\] worked out.

OpenStudy (mertsj):

\[-6x^3-10x^2-2x\]

OpenStudy (mertsj):

We have to subtract that from the first expression so we will just add the opposite:

OpenStudy (mertsj):

\[-6x^3-5x^2+18x+15+6x^3+10x^2+2x=5x^5+20x+15\]

OpenStudy (mertsj):

\[5x^2+20x+15=5(x^2+4x+3)=5(x+3)(x+1)\]

OpenStudy (mertsj):

So that's the numerator. Just put it over the denominator.

OpenStudy (s):

ooh, now i get that too, probably had a mistake while simplifying. Thank you.Now it makes sense!

OpenStudy (mertsj):

yw